A gigantic warehouse stores approximately 80 million empty aluminum beer and soda cans. Recently, a fire occurred at the warehouse. The smoke from the fire contaminated many of the cans with blackspot, rendering them unusable. A statistician was hired by the insurance company to estimate p, the true proportion of cans in the warehouse that were contaminated by the fire. How many aluminum cans should be randomly sampled to estimate p to within 0.08 with 95% confidence?
Answer #1

Solution :

Given that,

= 0.50

1 - = 0.50

margin of error = E = 0.08

Z/2 = 1.96

sample size = n = (Z / 2 / E)2 * * (1 - )

= (1.96 / 0.08)2 * (0.50 * 0.50)

= 150

sample size = n = 150